If r=40cm, F= 150N, and the torque has maximum allowed value of 8.2 N.m the angle between the force and the lever arm should be equal to ____
1
Expert's answer
2020-10-30T13:13:30-0400

By definition, the torque is the crossproduct of the force and the lever arm. Thus, its magnitude is given by:


"\tau = rF\sin\theta"

where "\theta" is the angle between the force and the lever arm. Expressing "\theta", obtain:


"\theta = \arcsin\left( \dfrac{\tau}{Fr} \right) = \arcsin\left( \dfrac{8.2N\cdot m}{150N\cdot 0.4m} \right) \approx 7.86\degree"

Answer. "7.86\degree".